if 15 workers do work in 4 days how much time required for 10 workers to complite same work
Answer by Edwin McCravy(20086)

Since 15 workers can do the job in 4 days, then twice as many or 30 workers can
do the job in half the time, 2 days. Therefore only 1/3 as many workers, that
is, 10 workers, would take 3 times as long or 6 days.

Or you can use the formula

Workers*days the first time = workers*days the second time

15·4 = 10·x
  60 = 10·x
   6 = x


Edwin